# Boosting Large-scale Parallel Training Efficiency with C4 : A Communication-Driven Approach

### v1

@@ 4.2. Results
The downtime in in June 2023 is broken down into four parts, according to the classification in Fig. 1. It reveals that the majority of the downtime was due to extended periods of system diagnosis. This was largely because identifying the precise cause of job failure could take hours or even days in the absence of our purpose-built diagnostic tools. The second major contributor to downtime was post-checkpoint overhead. That is because the users often scheduled infrequent checkpoints to save the model parameters, not anticipating high error rates in a large-scale training cluster. This led to a significant amount of computing activity being rendered useless between the last valid checkpoint and the moment the job failed due to an error. To address this issue and reduce the amount of wasted computing time, users have begun saving checkpoints more frequently by taking advantage of the high-performance infrastructure provided (Wang et al., 2023a). Besides the diagnosing and post-checkpoint cost, there is also 3.4% of time spent on crash detection. Typically, users are not immediately aware that their job has stalled following an error event. For instance, PyTorch jobs that might experience a hang for up to 30 minutes, until the elastic agent (Pytorch, [n. d.]) kills the processes. This delay creates a span of idle time for computing resources, and thus constitutes a non-negligible portion of the total downtime.
In December 2023, despite system diagnosis continuing to account for the majority of downtime, its contribution had been reduced by approximately 27x compared to past performance. which is somewhat lower than the average efficiency gains observed. The deployment of C4D has markedly accelerated both error detection and the pinpointing of faulty components, cutting down the response time to mere tens of seconds. Nevertheless, additional minutes are still required by the steering service to isolate the affected nodes and restart the job, indicating room for further improvement. The second leading cause of downtime remains the post-checkpoint duration. However, with the adoption of more frequent checkpointing, users can now save checkpoints every 10 minutes. As a result, the post-checkpoint downtime has been significantly reduced by 33x. The cost associated with job re-initialization has remained constant. It has decreased from 0.6% to 0.15% of the total time, because the average error rate has decreased by 3.33x, after the most vulnerable components of the system were identified and enhanced during this 6 months.
@@
4.2.2. The Evaluations on C4P’s Effectiveness.
Figure 8. Improvement from balanced traffic between bonded ports.
Balancing traffic between the two bonded physical ports. As discussed in Section 3.2, when flows are dispatched from two distinct physical ports constituting a single bonded port, there is a possibility that both flows could be routed to the same physical port on the receiving end, leading to an imbalance of traffic between the two physical ports in receivers. By employing C4P, we can designate a dedicated path for the flows in each physical port, thus preventing performance degradation caused by such imbalances. Running “nccltest”, the collective operation benchmarking tool supplied by NVIDIA, we are able to measure the average busbw, a metric that reflects the effective communication performance (higher busbw means lower communication delay). Fig. 8 shows the performance of a single allreduce operation with and without C4P enabled. Here, the results from the two-node (16-GPU) test case don’t align with the effective bandwidth observed in the network. This discrepancy can be attributed to the methodology employed to calculate busbw within the benchmarks. We will not delve into this particular issue in this paper, and instead concentrate on other test cases to better illustrate our research outcomes. Without C4P, the effective busbw is lower than 240 Gbps in most test cases, which is far from the ideal bandwidth of network (around 360 Gbps). However, with C4P activated, the effective busbw increases close to the peak value 360 Gbps, which means 50% performance gain, showcasing the tangible benefits of C4P in optimizing network performance.
@@
Figure 9. Effectiveness of global traffic engineering.
@@
Figure 10. The CNP count received in each bonded port.
It is important to clarify that the peak throughput in this system is constrained by the NVLink fabric, causing the maximum attainable bus bandwidth at 362 Gbps. Because of the limitation of NVLink fabric, the network’s capacity is underutilized, which results in an absence of queue buildup in switches and congestion-induced transmission rate reduction in senders. To evaluate the performance of C4P within a congested network environment, we intentionally reduced the number of active spine switches by half, effectively increasing the network oversubscription rate to 2:1. Under this new configuration, we replicated the previous experiments to observe how C4P handles increased traffic conditions. The outcomes of these tests are presented in Fig. 9b. A similar conclusion can be drawn from these outcomes, with the notable exception that the performance of concurrent tasks exhibits slight variations, even when C4P is enabled. Specifically, we observe a small gap of just 11.27 Gbps between the highest and lowest throughputs. This difference is attributed to the congestion control mechanism in RDMA NICs, which dynamically adjusts the data transmission rate at the sender side in response to network conditions. As depicted in Fig. 10, each bonded port receives approximately 15,000 Congestion Notification Packets (CNPs) every second, with the count fluctuating between 12,500 and 17,500. The arrival of CNPs instigates a throttling of the data transfer rate from the senders. Fluctuations in the frequency at which these CNPs are received directly contribute to the observed variability in the flows’ effective bandwidth. Despite these variations, the overall throughput is substantially improved by 65.55%, and the long-tail problem is obviously alleviated, with C4P enabled.
@@
Figure 11. Effectiveness of load balance when link failure appears.
Tolerance to dynamic link failures. The efficacy of global traffic engineering relies on the stability of the underlying network conditions. A downed link can necessitate the rerouting of impacted flows, which could lead to imbalanced traffic distribution within the network, and dismiss the efficacy of global traffic engineering. In this case, load balance mechanism would take into effect to adjust the load on each flow, thereby rebalancing network traffic. To assess the effectiveness of this mechanism, we replicated previous experiments in a 1:1 oversubscribed network and intentionally deactivated a link during the experiments. To obtain the tasks’ instant performance, other than averaged value after the tasks are finished, we refined our testing benchmark to output results immediately upon completion of each allreduce operation. And the evaluation results are shown in Fig. 11. Though high-frequency timestamp recording results in some fluctuations in the data, we can still observe significant improvement in system throughput, with dynamic load balance enabled. In the tests that C4P load balancing is not activated, the bus bandwidth for the given tasks experiences considerable degradation, with values ranging between 160 Gbps and 220 Gbps and an average bandwidth of 185.76 Gbps. Conversely, when C4P load balancing is enabled, the observed bus bandwidth improves significantly, ranging from 290 Gbps to 335 Gbps, with an average of 301.46 Gbps. This enhancement represents a substantial performance gain of 62.3% when a link error occurs during the training processes. Note that, with 1 link error among the 8 uplinks, the theoretical ideal performance would be 7/8 of the original, i.e., 315 Gbps. And the performance achieved with C4P enabled is found to be in close proximity to this ideal value.
@@
Figure 12. Comparison on switch port bandwidth with/without load balance.
@@
To demonstrate how C4P load balance improves the system throughput, we have collected the statistics from the leaf switches regarding the traffic distribution across each port, as illustrated in Fig. 12. With C4P traffic engineering implemented, all upstream ports exhibit near-optimal bandwidth utilization prior to the induction of a link error. However, the emergence of the link error leads to a marked discrepancy in throughput across different ports. In the absence of C4P load balancing, only three of the ports exhibit traffic increment, indicating that the flows intended for the disabled link are being rerouted to these ports. As a result, the available bandwidth for each flow in these congested ports is reduced, and the performance of the flows in the same communication channel is also negatively affected. Consequently, a noticeable reduction in bandwidth is observed in the other ports. Conversely, C4P load balancing dynamically adjusts the distribution of network traffic, decreasing the burden on congested flows while increasing it on those traversing underutilized paths. This strategy leads to a more balanced load across the ports, which allows those that were previously underutilized to handle more traffic. Thus, this optimization results in an overall enhancement of the system’s throughput.
@@
Performance improvement in real-life jobs. To evaluate the efficacy of C4P in enhancing the performance of real-world applications, we conducted tests using three representative training jobs: Job1 involves training a GPT model utilizing the Megatron framework, incorporating both TP and DP strategies for distributed training. This model comprises 22 billion parameters, and the TP and DP sizes are set to 8 and 16, respectively. Job2 trains a Llama model, which contains 7 billion parameters. For this task, the Deepspeed framework is employed for distributed training. The zero-optimization (Rajbhandari et al., 2020) is activated, with the training leveraging only DP. Job3 entails training another GPT model with 175 billion parameters. This model is also trained based on Megatron framework, using both TP and PP strategies. With both TP and PP sizes configured as 8, it effectively creates 2 DP groups.
Figure 13. Performance improvement in real-life jobs.
@@
The performance evaluation results are depicted in Fig. 13. As can be discerned from the figure, significant improvements in the performance of the first two jobs. Specifically, the throughput of Job1 has increased by 15.95%, rising from 74.82 to 86.76 samples per second. Similarly, Job2 shows an enhancement of 14.1%, with throughput climbing from 156.59 to 178.65 samples per second. In contrast, Job3 does not exhibit a noticeable improvement in performance. Our analysis suggests that the discrepancy in performance gains is closely associated with the proportion of communication time within a training step. For the first two jobs, the communication overhead constitutes over 30% of each iteration’s duration. However, Job3 is configured with a high Gradient Accumulation (GA) value of 16, which implies that parameter updates occur only once every 16 steps. This setting substantially reduces the relative communication cost, which explains the minimal performance improvement observed for Job3.
5. Related Works
@@
Improving stability by monitoring and fault tolerance. Existing research on data center monitoring primarily focuses on general applications within data centers (Yu et al., 2015; Joshi et al., 2018; Kutare et al., 2010), focusing on the behavior of hardware components such as CPUs, RDMA NICs, and switches. However, when compared to conventional data center applications, the training of LLMs is characteristically on a much larger scale and involves more complex interactions between software and hardware modules. This complexity necessitates more sophisticated, high-granularity monitoring solutions (Rasley et al., 2020; Nvidia, 2023b) that are integrated within the software stack.
@@
In terms of fault tolerance, the prevalent approach is the use of in-memory checkpointing (Wang et al., 2023b; Nicolae et al., 2020; Mohan et al., 2021; Eisenman et al., 2022). These efforts copy the current LLM states to the host memory with high frequency such as one checkpoint per iteration. In the event of an error, once the issue has been identified, training can resume from the most recent checkpoint.
@@
Improving performance by better training architectures and network. With given computation resources, the performance of LLM training can be improved by multiple-layer efforts, including (from the higher layer to the lower layer) training architectures, communication libraries, and network structure. The LLM training tries to parallelize GPU computation with communication and CPU operations (Zhang et al., 2023; Rasley et al., 2020; Shoeybi et al., 2019), or change the behavior of LLM training to reduce the overload, leveraging the properties of specific model structures (Cui et al., 2023; Liu et al., 2023). Communication libraries try to better utilize the given network bandwidth, including NVLink, PCIe, and RNIC, to shorten the delay of a data transmission (Jeaugey, 2017; Microsoft, 2023; Dong et al., 2021; Facebook, 2023), or design better strategies to aggregate data in collective communication (Shah et al., 2023; Cowan et al., 2023; Cai et al., 2021). Network structure research explore reasonable topologies to connect multiple hosts with appropriate bandwidth, providing better performance in LLM training.
6. discussion
Adaptive routing (Dally and Seitz, 1987) and packet spraying (Dixit et al., 2013) generally point to the same solution, which involves performing per-packet load balancing within switches based on network conditions. This approach holds advantages over traditional ECMP technology as it can effectively address imbalances in network bandwidth utilization caused by large flow hash collisions. However, implementing such techniques introduces complexities for RoCE networks and higher-level applications. For instance, to alleviate the pressure on RNIC’s on-chip buffers due to out-of-order packets, NVIDIA’s RNIC only enable adaptive routing for RDMA Write operations (Nvidia, 2023a). Consequently, all RDMA Write packets are transformed into RDMA Write Only packets, ensuring each packet carries the destination DMA address. Furthermore, to preserve the ordering required by upper-layer applications for RDMA operations, the final packet must use the RDMA Write with immediate operation, and it can only be sent after all preceding packets have been acknowledged. This safeguards the integrity of the Completion Queue Entry (CQE) generation, ensuring that all data from Write operations is properly committed to memory. These constraints necessitate careful adaptation of upper-layer software to enable adaptive routing. Currently, AR support has been integrated into NCCL, and has been successfully verified on InfiniBand networks. Despite our active exploration of this technology, it has not achieved widespread adoption, primarily due to two significant reasons. First, AR on RoCE depends on underlying hardware and requires the latest switches and RNICs for support. Second, AR technology relies on a lossless RDMA network, which can exacerbate congestion when endpoint congestion occurs, potentially paralyzing the entire cluster network (Fu and Kim, 2017). Actually, to avoid the operational risks associated with lossless networks, we have extensively deployed lossy RDMA networks. Achieving optimal AR performance in lossy RDMA networks is also a problem we are currently exploring.
7. Conclusion

### v2

@@ IV-B Results
The downtime in in June 2023 is broken down into four parts, according to the classification in Fig. 2. It reveals that the majority of the downtime was due to extended periods of system diagnosis. This was largely because identifying the precise cause of job failure could take hours or even days in the absence of our diagnostic tools. The second major contributor to downtime was post-checkpoint overhead. That is because the users often scheduled infrequent checkpoints to save the model parameters, not anticipating high error rates in a large-scale training cluster. This led to a significant amount of computing activity being rendered useless between the last valid checkpoint and the moment the job failed due to an error. To address this issue and reduce the amount of wasted computing time, users have begun saving checkpoints more frequently by taking advantage of the high-performance infrastructure provided [53]. Besides the diagnosing and post-checkpoint cost, there is also 3.4% of time spent on crash detection. Typically, users are not immediately aware that their job has stalled following an error event. For instance, PyTorch jobs that might experience a hang for up to 30 minutes, until the elastic agent [42] kills the processes. This delay creates a span of idle time for computing resources, and thus constitutes a non-negligible portion of the total downtime.
In December 2023, despite system diagnosis continuing to account for the majority of downtime, its contribution had been reduced by approximately 27x compared to past performance. which is somewhat lower than the average efficiency gains observed. The deployment of C4D has markedly accelerated both error detection and the pinpointing of faulty components, cutting down the response time to mere tens of seconds. Nevertheless, additional minutes are still required by the steering service to isolate the affected nodes and restart the job, indicating room for further improvement. The second leading cause of downtime remains the post-checkpoint duration. However, with the adoption of more frequent checkpointing, users can now save checkpoints every 10 minutes. As a result, the post-checkpoint downtime has been significantly reduced by 33x. The cost associated with job re-initialization has remained constant. It has decreased from 0.6% to 0.15% of the total time, because the average error rate has decreased by 3.33x, after the most vulnerable components of the system were identified and enhanced during this six months.
@@
IV-B2 The Evaluations on C4P’s Effectiveness.
Figure 9: Improvement from balanced traffic between bonded ports.
Balancing traffic between the two bonded physical ports. As discussed in Section II-D, when two flows are dispatched from two distinct physical ports of a NIC, there is a possibility that both flows could be routed to the same physical port on the receiving end, leading to an imbalance of traffic between the two physical ports in receivers. By employing C4P, we can designate a dedicated path for the flows in each physical port, thus preventing performance degradation caused by such imbalances. Running “nccltest”, the collective operation benchmarking tool supplied by NVIDIA, we are able to measure the average busbw, a metric that reflects the effective communication performance (higher busbw means lower communication delay). Fig. 9 shows the performance of a single allreduce operation with and without C4P enabled. Here, the results from the two-node (16-GPU) test case don’t align with the effective bandwidth observed in the network. This discrepancy can be attributed to the methodology employed to calculate busbw within the benchmarks. We will not delve into this particular issue in this paper, and instead concentrate on other test cases to better illustrate our research outcomes. Without C4P, the effective busbw is lower than 240 Gbps in most test cases, which is far from the ideal bandwidth of network (around 360 Gbps). However, with C4P activated, the effective busbw increases close to the peak value 360 Gbps, which means 50% performance gain, showcasing the tangible benefits of C4P in optimizing network performance.
@@
Figure 10: Effectiveness of global traffic engineering.
@@
Figure 11: The CNP count received in each bonded port.
It is important to clarify that the peak throughput in this system is constrained by the NVLink fabric, causing the maximum attainable bus bandwidth at 362 Gbps. Because of the limitation of NVLink fabric, the network’s capacity is underutilized, which results in an absence of queue buildup in switches and congestion-induced transmission rate reduction in senders. To evaluate the performance of C4P within a congested network environment, we intentionally reduced the number of active spine switches by half, effectively increasing the network oversubscription rate to 2:1. Under this new configuration, we replicated the previous experiments to observe how C4P handles increased traffic conditions. The outcomes of these tests are presented in Fig. 10b. A similar conclusion can be drawn from these outcomes, with the exception that the performance of concurrent tasks exhibits slight variations, when C4P is enabled. Specifically, we observe a small gap of just 11.27 Gbps between the highest and lowest throughputs. This difference is attributed to the congestion control mechanism in RDMA NICs, which dynamically adjusts the data transmission rate at the sender side in response to network conditions. As depicted in Fig. 11, each bonded port receives approximately 15,000 Congestion Notification Packets (CNPs) every second, with the count fluctuating between 12,500 and 17,500. The arrival of CNPs instigates a throttling of the data transfer rate from the senders. Fluctuations in CNP receiving frequency directly contribute to the observed variability in the flows’ effective bandwidth. Despite these variations, the overall throughput is substantially improved by 65.55%, and the long-tail problem is obviously alleviated.
@@
Figure 12: Effectiveness of load balance when link failure appears.
Tolerance to dynamic link failures. The efficacy of global traffic engineering relies on the stability of the underlying network conditions. A down-link can necessitate the rerouting of impacted flows, which could lead to imbalanced traffic distribution within the network, and dismiss the efficacy of global traffic engineering. In this case, load balance mechanism would take into effect to adjust the load on each flow, thereby rebalancing network traffic. To assess the effectiveness of this mechanism, we replicated previous experiments in a 1:1 oversubscribed network and intentionally deactivated a link during the experiments. To obtain the tasks’ instant performance, other than averaged value after the tasks are finished, we refined our testing benchmark to output results immediately upon completion of each allreduce operation. And the results are shown in Fig. 12. Though high-frequency timestamp recording results in some fluctuations in the data, we can still observe significant improvement in system throughput, with dynamic load balance enabled. In the tests that C4P load balancing is not activated, the bus bandwidth for the given tasks experiences considerable degradation, with values ranging between 160 Gbps and 220 Gbps and an average bandwidth of 185.76 Gbps. Conversely, when C4P load balancing is enabled, the observed bus bandwidth improves significantly, ranging from 290 Gbps to 335 Gbps, with an average of 301.46 Gbps. This enhancement represents a substantial performance gain of 62.3% when a link error occurs during the training processes. Note that, with 1 link error among the 8 uplinks, the theoretical ideal performance would be 7/8 of the original, i.e., 315 Gbps. And the performance achieved with C4P enabled is found to be in close proximity to this ideal value.
@@
Figure 13: Comparison on switch port bandwidth with/without load balance.
@@
To demonstrate how C4P load balance improves the system throughput, we have collected the statistics from the leaf switches regarding the traffic distribution across each port, as illustrated in Fig. 13. With C4P traffic engineering implemented, all upstream ports exhibit near-optimal bandwidth utilization prior to the induction of a link error. However, the emergence of the link error leads to a marked discrepancy in throughput across different ports. In the absence of C4P load balancing, only three of the ports exhibit traffic increment, indicating that the flows intended for the disabled link are being rerouted to these ports. As a result, the available bandwidth for each flow in these congested ports is reduced, and the performance of the flows in the same communication channel is also negatively affected. Consequently, a noticeable reduction in bandwidth is observed in the other ports. Conversely, C4P load balancing dynamically adjusts the distribution of network traffic, decreasing the burden on congested flows while increasing it on those underutilized paths. This strategy leads to a more balanced load across the healthy links, and results in an overall enhancement of the system’s throughput.
@@
Performance improvement in real-life jobs. To evaluate the efficacy of C4P in enhancing the performance of real-world applications, we conducted tests using three representative training jobs: Job1 involves training a GPT model utilizing the Megatron framework, incorporating both TP and DP strategies for distributed training. This model comprises 22 billion parameters, and the TP and DP sizes are set to 8 and 16, respectively. Job2 trains a Llama model, which contains 7 billion parameters. For this task, the Deepspeed framework is employed for distributed training. The zero-optimization [45] is activated, with the training leveraging only DP. Job3 entails training another GPT model with 175 billion parameters. This model is also trained based on Megatron framework, using both TP and PP strategies. With both TP and PP sizes configured as 8, it effectively creates 2 DP groups.
Figure 14: Performance improvement in real-life jobs.
@@
The performance evaluation results are depicted in Fig. 14. As can be discerned from the figure, significant improvements in the performance of the first two jobs. Specifically, the throughput of Job1 has increased by 15.95%, rising from 74.82 to 86.76 samples per second. Similarly, Job2 shows an enhancement of 14.1%, with throughput climbing from 156.59 to 178.65 samples per second. In contrast, Job3 does not exhibit a noticeable improvement in performance. Our analysis suggests that the discrepancy in performance gains is closely associated with the proportion of communication time within a training step. For the first two jobs, the communication overhead constitutes over 30% of each iteration’s duration. However, Job3 is configured with a high Gradient Accumulation (GA) value of 16, which implies the communication cost is reduced by 16x. This setting substantially reduces the relative communication cost, which explains the minimal performance improvement observed for Job3.
V Related Work and Discussion
@@
Improving stability by monitoring and fault tolerance. Existing research on data center monitoring primarily focuses on general applications within data centers [57, 23, 25], focusing on the behavior of hardware components such as CPUs, RDMA NICs, and switches. However, when compared to conventional data center applications, the training of LLMs is characteristically on a much larger scale and involves more complex interactions between software and hardware modules. This complexity necessitates more sophisticated, high-granularity monitoring solutions [46, 40] that are integrated within the software stack.
@@
In terms of fault tolerance, the prevalent approach is the use of in-memory checkpointing [54, 34, 31, 16]. These efforts copy the current LLM states to the host memory with high frequency such as one checkpoint per iteration. In the event of an error, once the issue has been identified, training can resume from the most recent checkpoint.
In addressing the challenge of straggler components, Jiang et al. [22] proposed an optimization solution to enhance the efficiency of large-scale AI clusters. Specifically, they identified slow components by monitoring the execution time of specific computing kernels. However, this approach is not suitable for cloud environments.
@@
Improving performance by better training architectures and network. With given computation resources, the performance of LLM training can be improved by multiple-layer efforts, including (from the higher layer to the lower layer) training architectures, communication libraries, and network structure. The LLM training tries to parallelize GPU computation with communication and CPU operations [58, 46, 48], or change the behavior of LLM training to reduce the overload, leveraging the properties of specific model structures [9, 28]. Communication libraries try to better utilize the given network bandwidth, including NVLink, PCIe, and RNIC, to shorten the delay of a data transmission [21, 30, 14, 17], or design better strategies to aggregate data in collective communication [47, 8, 6]. Network structure research explore reasonable topologies to connect multiple hosts with appropriate bandwidth, providing better performance in LLM training.
Adaptive routing [10] and packet spraying [12] generally point to the same solution, which involves performing per-packet load balancing within switches based on network conditions. The decision to implement adaptive routing is entirely up to the end users, who must configure the related condition variables of CCL. However, this approach poses significant challenges, as the efficiency of adaptive routing can be compromised by the flows that are deterministic routed.
Applicability and Limitations of C4. C4D has broad applicability in business AI training tasks that rely on collective communication. Although we demonstrate its effectiveness using DP as an example, it is natural to wonder whether it can also be used in other scenarios, such as Pipeline Parallel (PP) [20] and Expert Parallel (EP) [26]. While C4D may not be able to detect non-communication hang or slow events within PP stages, as it relies on send/recv operations rather than collective communication, any resulting hang or slow effects would propagate to subsequent DP stages and could still be caught by C4D. In the case of EP, load imbalance among workers may occur, which can be mitigated by averaging collected data over a predefined period to smooth out random variations and highlight systemic issues. In future work, we plan to incorporate load variation into C4D to enable more precise slow detection.
The primary limitations of C4 includes: C4D’s error detection capabilities rely on syndromes observed through our enhanced communication library. Consequently, C4D has limited ability to detect errors occurring during the job initialization phase, before collective communication operations are initiated. And deploying C4P requires detailed knowledge of the system configuration, including network topology, the number of switches at each layer, the number of ports per switch. Adapting C4P to different clusters necessitates retrieving this information from a background management system.
VI Acknowledgment
This work was partially supported by Research Grants Council of HKSAR (16213824). We thank Yu Zhou for developing the initial version of C4D.
VII Conclusion